HETAV PATEL

Grade 11,

1) if A+B+C=3.14, then prove that, cosA+cosB+cosC 2) if A+B+C=3.14,then prove that tan^2A/2+tan^2B/2+tan^2C/2>1 sir please solve these two question by using graph

1) if A+B+C=3.14, then prove that, cosA+cosB+cosC<3/2, where A,B,c are distinct.

2) if A+B+C=3.14,then prove that tan^2A/2+tan^2B/2+tan^2C/2>1

sir please solve these two question by using graph

Grade:11

1 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

Using CosA = [ (b2 + c2 – a2) / 2bc ]  etc. , in the relation  cosA  +  cosB  +  cosC  =  3/2  we obtain


[ (b2 + c2 – a2) / 2bc ] + [ (c2 + a2 – b2) / 2ac ] + [ (a2 + b2 – c2) / 2ab ] = 3/2

 

or   a.(b2 + c2 – a2)  +  b.(c2 + a2 – b2)  + c.(a2 + b2 – c2) = 3abc

 

or   a.(b2 + c2)  +  b.(c2 + a2 )  + c.(a2 + b2 ) =  a3 + b3 + c3 + 3 abc

 

or   ab2 + ac2 +  bc2 + ba2   + ca2 + cb2 =  a3 + b3 + c3 + 3 abc



Now subtract 6 abc from both the sides and rearraginf the terms we obtain



ab2 + ac2 - 2abc +  bc2 + ba2  - 2abc  + ca2 + cb2  - 2abc = a3 + b3 + c3 - 3 abc

 

or  a ( b - c )2 + b ( c – a )2 + c ( a – b )2  =  a3 + b3 + c3 - 3 abc

 

or  a ( b - c )2 + b ( c – a )2 + c ( a – b )2  =  (a + b + c) (a2 + b2 + c2 – ab – bc – ca)                    ... ( 1 )



Multiply both the sides by 2 we get



Left hand side = 2a ( b - c )2 + 2b ( c – a )2 + 2c ( a – b )2



&   Right hand side = (a + b + c) (2a2 + 2b2 +2 c2 – 2ab – 2bc – 2ca) 

 

or RHS = (a + b + c)  (a2 + b2 + c2 – 2ab – 2bc – 2ca + a2 + b2 + c2

 

 = ( a + b + c ) [ (a2 + b2 – 2ab) + ( b2 + c2 – 2 bc ) + (a2 + c2 - 2 ca) ]

 

 = ( a + b + c ) [ ( a – b )2 + (b – c )2 + ( c – a )2 ]



Now eqn (1) becomes

 

2a ( b - c )2 + 2b ( c – a )2 + 2c ( a – b )2 =  ( a + b + c ) [ ( a – b )2 + (b – c )2 + ( c – a )2 ]

 

or (a - b - c) ( b - c )2 + (b -a - c) ( c – a )2 + ( c- a - b) ( a – b )2 = 0       ... (2)


Now we know that sum of two sides of a triangle is greater than the third side. that is  (a - b - c) is not equal to 0 similarly (b - a - c) is not equal to 0 and also (c - a - b) is not equal to 0



Hence in eqn (2) squared terms must be equal to zero.



ie. ( b - c )2 = 0  or  b = c . Similarly c = a



Hence a = b = c

 

 i.e. The triangle is Equilateral.

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi